# Token-free LLMs Can Generate Chinese Classical Poetry with More Accurate Format

### 07 Jan 2024 v1

@@ Abstract.
We then show that a token-based model can be easily tailored into a token-free model (in terms of Chinese), which can largely solve the format accuracy problem. Our tailoring procedure removes long-token from vocabulary and keeps only character-level or byte-level tokens.
As part of our contribution, we release the finetuned token-free model (which is based on Qwen-chat-7B), which can generate chinese classical poetry following complex instructions like LLMs (such as story paraphrasing), and also perform well in format. On the test set, our token-free model achives an format accuracy of 0.96, compared to 0.84 for token-based counterparts and 0.38 for GPT-4.
@@ 1. Introduction
In this example, GPT-4 is asked to write a poem the form of ”Rumengling”, with the key word ”cheerful”. We can see that the generated poem is relative in content, but clearly violate the format requirement, even with an exmaple format provided in the prompt. To present it clearer to the readers, the redundant characters are marked with a strikethrough in the above example.
Prompt 请写一首如梦令,关键词为“兴高采烈”。请严格按照如梦令对每一句话的字数要求,下面给出一个例子: 常记溪亭日暮,沉醉不知归路。 兴尽晚回舟,误入藕花深处。 争渡,争渡,惊起一滩鸥鹭。Äpgjy Response 浅笑嫣然花开时,兴高采烈醉春风。 半醺半醉半睡眠,悠然步入花丛中。 漫步,漫步,蝴蝶舞动春光浓。
@@ 2.1. Comparison with existing general LLMs
2.2. Comparison with By-T5, Jiuge-GPT2 and other poetry generating models
@@ 3.1. Method
Prompt List all the charaters in the following token: $<|extra\_1|>$大模型$<|extra\_1|>$Äpgjy Response $<|extra\_1|>$$<|extra\_1|>$$<|extra\_1|>$$<|extra\_1|>$
@@ 4. Experiment2: Can token-free models do better?
Since the token-based models have been proved to be defective in token-character knowledge, one of the solutions is to use a token-free model. In this section, we will demonstrate how we tailor existing token-based LLMS into token-free models and preserves its pretrained ability to some extent. And we will show how this tailored token-free model could enhance the format accuracy in poetry generation.
@@ 4.1. Method
Task design. We hope the model could deal with complex task such as story paraphrasing. The final user interface of our model is shown in Figure 2. User instructions are composed of a format requirement and a content requirement.
Format requirement. Any format type of SHI or CI could be chosen. In the above demonstration, we provide only 12 format types for interface simplicity.
@@ 4.2. Results
User: 简要介绍一下日心说Äpgjy Agent 日心说,也称日心地动说,是16th century末期由哥白尼提出的天体学说,它将日、月、水星、金星、木星、土星、天王星、海王星和冥王星等9大天体,及小天体如彗星、小星体等,都绕着以日为圆心的椭圆轨转。
User: Briefly introduce the heliocentric theory.Äpgjy Agent The heliocentric theory is a model of the solar system that places the Sun at the center, with the planets orbiting around it in椭ical paths. This theory was proposed by Nicolaus Copernicus in the 16th century, but it was not widely accepted until the work of later astronomers such as Johannes Kepler and Galileo Galilei. The heliocentric model is now the most widely accepted model of the solar system, and it is supported by a vast amount of observational evidence.
@@
Content Relevance. An example prompt comparable to the GPT-4 case is provided below. Here the model is required to write a poem in the form of ”Rumengling” and with the keyword ”cheerful”. In the example, the model generates a poem translated as ”Speaking of the autumn weather, sharing thoughts with others, in high spirits and delight. Seek no more, seek no more. Unaware when the wine arrives”.
@@ 5. Conclusion
We train our token-free model on the poetry generation instructions dataset. Our model is based on Qwen-chat-7B. The tailored model achives an format accuracy of 0.96, much higher than its token-based counterpart (0.84), and also higher than GPT-4 (0.38).

### 09 Jan 2024 v2

@@ Abstract.
We then show that a token-based model can be easily tailored into a token-free model (in terms of Chinese), which can largely solve the format accuracy problem. Our tailoring procedure removes long-tokens from the vocabulary and the language model head, and keeps only character-level or byte-level tokens.
As part of our contribution, we release the finetuned token-free model (which is based on Qwen-chat-7B), which can generate chinese classical poetry following complex instructions like LLMs (such as story paraphrasing), and also perform well in format. On the test set, our token-free model achives an format accuracy of 0.96, compared to 0.84 for token-based equivalents and 0.38 for GPT-4.
@@ 1. Introduction
In this example, GPT-4 is asked to write a poem the form of ”Rumengling”, with the key word ”cheerful”. We can see that the generated poem is relative in content, but clearly violate the format requirement, even with an exmaple format provided in the prompt. To present it clearer to the readers, the redundant characters are marked with a strikethrough.
Prompt 请写一首如梦令,关键词为“兴高采烈”。请严格按照如梦令对每一句话的字数要求,下面给出一个例子: 常记溪亭日暮,沉醉不知归路。 兴尽晚回舟,误入藕花深处。 争渡,争渡,惊起一滩鸥鹭。 Äpgjy Response 浅笑嫣然花开,兴高采烈醉春 半醺半醉半睡眠,悠然步入花丛 漫步,漫步,蝴蝶舞动春光
@@ 2.1. Comparison with existing general LLMs
2.2. Comparison with ByGPT5, Jiuge-GPT2 and other poetry generating models
@@ 3.1. Method
Prompt List all the charaters in the following token: $<|extra\_1|>$大模型$<|extra\_1|>$pgjy Response $<|extra\_1|>$$<|extra\_1|>$$<|extra\_1|>$ $<|extra\_1|>$
@@ 4. Experiment2: Can token-free models do better?
Since the token-based models have been proved to be defective in token-character knowledge, one of the solutions is to use a token-free model. In this section, we will demonstrate how we tailor existing token-based LLMS into token-free models and preserves its pretrained ability to some extent. And we will show how this tailored token-free model could enhance the format accuracy in poetry generation. The detailed token-free model framework is shown in Figure 1.
@@ 4.1. Method
Task design. We hope the model could deal with complex task such as story paraphrasing. The final user interface of our model is shown in Figure 2.
User instructions are composed of a format requirement and a content requirement.
Format requirement. Any format type of SHI or CI could be choosen. In the demonstration of our AI Poet (Figure 2), we provide only 12 format types for interface simplicity.
@@ 4.2. Results
User: 简要介绍一下日心说Äpgjy Agent: 日心说,也称日心地动说,是16th century末期由哥白尼提出的天体学说,它将日、月、水星、金星、木星、土星、天王星、海王星和冥王星等9大天体,及小天体如彗星、小星体等,都绕着以日为圆心的椭圆轨转。
User: Briefly introduce the heliocentric theory.Äpgjy Agent: The heliocentric theory is a model of the solar system that places the Sun at the center, with the planets orbiting around it in椭ical paths. This theory was proposed by Nicolaus Copernicus in the 16th century, but it was not widely accepted until the work of later astronomers such as Johannes Kepler and Galileo Galilei. The heliocentric model is now the most widely accepted model of the solar system, and it is supported by a vast amount of observational evidence.
@@
Content Relevance. An example model output comparable to the GPT-4 case in the Introduction section is provided below. Here the model is required to write a poem in the form of ”Rumengling” and with the keyword ”cheerful”. In the example, the model generates a poem translated as ”Talking about the autumn weather, sharing thoughts with others, in high spirits and delight. Ask no more, ask no more. Drinking too much, losing consciousness, and not knowing where I am.”
Response 笑口频开深院,更说秋风天气。 心事向人知,却好兴高采烈。 休觅,休觅,酒到不知醒地。
@@ 5. Conclusion
We train our token-free model on the poetry generation instructions dataset. Our model is based on Qwen-chat-7B. The tailored model achives an format accuracy of 0.96, much higher than its token-based equivalent (0.84), and also higher than GPT-4 (0.38).